OpenStudy (anonymous):

The length of a rectangle is 2 more than 3 times the width. If the perimeter is 100 meters, what is the width of the rectangle? 13 meters 12 meters 11 meters

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

l - length w - width perimeter = 2l + 2w l = 3w + 2 perimeter = 100 100 = 2l + 2w So..what's the first step?

OpenStudy (anonymous):

I'm confused, what?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

You have these two equations: 100 = 2l + 2w l = 3w + 2

OpenStudy (studygurl14):

do you get how i got them?

OpenStudy (anonymous):

oh ok I already got the answer

OpenStudy (anonymous):

took some time to process lol

OpenStudy (anonymous):

nvm thx

OpenStudy (studygurl14):

:)